# The second data release of the Indian Pulsar Timing Array: Investigating the Effect of Coronal Mass Ejection on PSR J1022+1001 and a Possible Mode Change of PSR J2145-0750

### v1

@@ 3.3 In situ data
The in situ data from STEREO-A reveals that the cme1, after starting from the Sun at 06-08-2022 01:48, with a speed of 900 km/s reaches STEREO-A around 08-08-2022 00:00 UTC; we see an enhancement around that time in the measured speed and temperature of the solar wind (see Fig. 4 - generated using CDAWeb [^4] Coordinated Data Analysis Web (CDAWeb) is part of NASA’s Space Physics Data Facility (SPDF), which serves as a central archive and access point for heliophysics and space weather data from various NASA and international missions - https://cdaweb.gsfc.nasa.gov/). The full data coverage is however missing due to some unknown reason, which could be the reason we do not see the expected jump in the plasma density. Nevertheless, we see subsequent increases in the plasma density after 10-08-2022 00:00 UTC, which notably corresponds to the cme2, which after starting from the Sun at 07-08-2022 11:00 UTC, with a speed of 776 km/s is supposed to reach STEREO-A around 09-08-2022 17:00 UTC. The differences in the actual arrival time of the CME and the calculated one lie in the simplistic assumptions, such as constant velocity, which is not completely valid given the fact that CME can accelerate or decelerate along its path due to the drag force [53, 54, 55]. The second CME (cme2) seems to be a possible secondary CME coming from nearly the same active region (as cme1, see Table 1), that got accelerated by possible interaction with the fast solar wind component.
@@ 3.4 DM excess estimate
Here, we estimate the DM-excess from the in situ data and the spherically symmetric SW model. From a spherically symmetric SW model consideration, the formula for the DM contribution coming from a portion of the LOS, say $P_{1}$-$P_{2}$ (see Fig. 5(a)) is
$DM_{P_{1}-P_{2}}=\frac{n_{sat}(r_{sat}(\text{AU}))^{2}}{1\penalty 10000\ \text{AU}}\int_{\theta 1}^{\theta 2}d\theta,$ (3.1)
where $n_{sat}$ is the electron density at the position of the satellite of interest, $r_{sat}\penalty 10000\ (\text{AU})$ is the distance of the satellite from the Sun in units of AU, $\theta_{i}$ is the angle subtended by the sun-$P_{i}$ line on the earth-pulsar LOS, where $i\in\{1,2\}$.
@@
Using the position coordinates of the satellite of interest, the pulsar’s LOS, and the properties of the CME modelling as per DONKI, we identify the portion of the LOS that gets influenced by the cme2’s impact on the day of our CME-affected-observation, see Fig. 5(b).
The $n_{sat}$ from the in situ data of STEREO-A is roughly 10 cm-3. Integrating along the affected portion of the LOS, in Eq. 3.1, we get the DM due to both the CME and the background SW $\sim$ $5.7\times 10^{-3}$ pc cm-3. From the in situ data over a long time baseline, the background electron density on average is $\sim$ 5 cm-3. From this, the DM excess expected due to the background SW is $2.8\times 10^{-3}$ pc cm-3. Hence the DM-excess from the in situ data is $\sim$ $2.9\times 10^{-3}$ pc/cc. This value conforms with the excess DM observed for PSR J1022+1001 on 09 August 2022 (see Section 3.1).
@@ 4 Analyzing a possible mode change epoch of PSR J2145-0750
PSR J2145-0750 is one of the brightest pulsars in the InPTA sample, with a distinct multi-component profile morphology, making it a natural candidate to investigate profile stability in detail. The DM time series of PSR J2145-0750 shows a lot of small-time-scale deviations around its median value, with a prominent outlier lying well below the median belt (see Fig. 6). Interestingly, not all of these deviations are due to poor detection – there is a significant number of such data points showing good detection with high S/N, indicating the possibility of profile showing marginal variability w.r.t. its template. To examine further, we proceeded as follows.
Figure 6: DM time series of PSR J2145-0750 along the InPTA-DR2 baseline (all prominent outliers, which are artifacts of poor data, are not included in this plot), highlighting the single outlier, corresponding to a potential mode-change, with red circle.
@@ 4.1 Peak-ratio distribution
For each observational epoch, we updated the corresponding archive/fits files using the InPTA-DR2 final timing solution. The header DM values were then corrected with epoch-wise estimates obtained from DMCalc, and the profiles were subsequently dedispersed to remove dispersion-induced artifacts. For ready comparison and consistent analysis of all the profiles across the baseline, we positioned them at the same pulse-phase (kept the main peaks at the center). In addition, the profiles were band-equalized [40] so that the power remained uniform across the frequency channels.
We then quantified the mode of the pulse profile by evaluating the peak ratio, defined as the ratio of the left peak to the right peak in the two prominent components of the profile, for each epoch. The resulting distribution of the ratio of the component amplitudes is well described by a Gaussian with mean and mode value of 1.0698 and a standard deviation of 0.0486, yielding 1-$\sigma$ bounds of [1.0212, 1.1183] (see Fig. 7). However, we identify a singular ‘outlier’, lying outside the Gaussian envelope, with a peak ratio of $\sim$0.8, observed at MJD 59468 (11 September 2021), the same outlier seen in the DM time series plot in Fig. 6, indicative of a possible mode change.
@@ 4.3 Eliminating systematics degenerate with mode change
The peak-ratio distribution reveals a single stable mode, well described by a Gaussian, and a distinct outlier corresponding to the epoch MJD 59468. The latter lies outside the Gaussian envelope and therefore raises the natural question of whether this deviation represents a genuine mode-change event or merely an artifact arising from observational systematics. Potential sources of such artifacts include: inter-stellar dispersion, which can alter the pulse shape; inter-stellar scintillation, which modulates the pulsar signal across frequency channels and can affect the frequency-integrated profile (particularly relevant for J2145–0750, which exhibits strong frequency-dependent profile variations); polarization calibration errors or flips in polarization channels; and radiometer noise, which may distort the band shape or modify the root-mean-square (RMS) of the profile baseline.
Each of these systematics has been carefully accounted for as follows. Dispersion effects are eliminated by using highly precise, epoch-wise DM estimates obtained with DMCalc on dual-band observations. Scintillation is accounted for by inspecting subbanded profiles for the four epochs of interest, which averages out the variation in signal strength across frequency channels. Band-shape equalization and phase bin collapsing suppress the impact of radiometer noise by reducing the RMS of the noise floor. The polarization calibration related systematics would be ineffective in our dataset since they are total intensity data rather than the full four Stokes parameter recordings. It is for this reason, however, an intrinsic profile change in polarized components would remain undetected in our data. Additionally, PSR J2145-0750 was at a separation angle of $\sim 160^{\circ}$ with the Sun, at the potential mode-changing epoch MJD 59468, thereby ruling out any solar origin for this outlier. Furthermore, we checked the observatory settings and the observatory logs for that epoch MJD 59468 – we found no mention of any observatory related glitches or atmospheric events, that could have attributed to such a potential mode change; also the observatory settings for that day was the standard one used for InPTA observations.
@@ 4.4 Indication of potential intrinsic magnetospheric changes
Apart from the change in the relative heights of the peaks of the profile we also see a symmetric broadening of the pulse width. We explore this, as a final test, to establish the plausible theoretical reason behind the potential mode change. We convolve each of the subbands of the template with a symmetric Gaussian kernel and find the best fit (using $\chi^{2}$ test) which resembles the actual subbanded profile. The standard deviation ($\sigma$) of the best-fit Gaussian kernel [^5] The parameters for the Gaussian convolution is $\{A,B,s,\sigma\}$, where A is the amplitude, B is the base line offset, s is the phase-shift. provides an estimate of the observed broadening in the corresponding subbanded profile. From the results of this analysis (see Fig. 10), we find non-zero $\sigma$ (in bin units) for each of the subbands.
@@
(b) subband 2 (central freq = 461 MHz)
(c) subband 3 (central freq = 436 MHz)
(d) subband 4 (central freq = 411 MHz)
(e) subband 5 (central freq = 386 MHz)
(f) subband 6 (central freq = 361 MHz)
(g) subband 7 (central freq = 336 MHz)
(h) subband 8 (central freq = 311 MHz)
@@
$\Delta t_{\Delta DM}\penalty 10000\ (ms)=8.3\times 10^{6}\times\Delta DM\times\frac{\Delta\nu}{\nu^{3}}\penalty 10000\ =\penalty 10000\ \sigma\penalty 10000\ (ms)=\sigma\penalty 10000\ (bins)\times\frac{P}{nbin}\penalty 10000\ ,$ (4.1)
where $P$ is the pulsar spin period, $\Delta\nu$ is the width of the subband, with $\nu$ being the central frequency of the subband. We observe that the estimated $\Delta DM\sim 0.08\penalty 10000\ pc/cc$, for all the subbands, exceeds the uncertainty $\sim 0.001\penalty 10000\ pc/cc$ in DM evaluation using DMCalc, by nearly two orders of magnitude. Hence, such a huge error666here by DM-error we refer to the difference in true DM and DMCalc-evaluated DM, while by DM-uncertainty we refer to the confidence interval on our DMCalc-evaluated DM value. in DM would have been clearly visible as a parabolic drift in the frequency-phase plot of the dedispersed frequency-integrated profile, which we do not see. Therefore, the source for the observed pulse-width broadening is less likely to be due to dispersion effects. Scattering effect, if any, is also possibly not the dominating contributor to such a pulse-width broadening, as validated from the absence of well-defined asymmetric exponential tails of the individual pulses in the profile and that clear scintles of 3–4 MHz wide scintillation bandwidths were seen in the observations. The symmetric broadening in the pulse-widths, therefore, in all possibilities, indicate a potential magnetospheric origin [56, 57].
@@ 5 Conclusions
In this work, we investigated astrophysically motivated outliers in the InPTA-DR2 DM time series, focusing on two representative pulsars: PSR J1022+1001 and PSR J2145-0750. We demonstrated that such outliers can serve as powerful probes of diverse plasma processes along the line of sight as well as intrinsic changes in a pulsar.
For PSR J1022+1001, we identified a distinct DM excursion on MJD 59800 (09 August 2022), which is geometrically consistent with a coronal mass ejection (CME) event detected by LASCO/C3 coronagraph, STEREO-A satellite and the DONKI database. Quantitative comparison between the DM excess inferred from radio data and that derived from in situ solar plasma measurements revealed excellent agreement, strongly establishing one of the first direct radio signatures of a CME along this particular PTA pulsar’s line of sight. This result underscores the potential for tracing heliospheric plasma dynamics and properties from the PTA datasets at astronomical distances where satellites are not present. Additionally, the number of IPTA pulsars monitored all across the sky is large enough, compared to space-based satellites, to possibly generate a three dimensional understanding of the outflow and interactions of solar wind and CMEs. The combination of several PTA datasets, across larger baselines, has the potential to contain more such solar-activity-induced outliers - this present understanding will therefore help in better modelling of the red-noise in the PTA dataset, induced by solar activities. This will lead to more accurate timing without having to throw away a significant portion of the dataset, at low separation angles, in the absence of better solar wind modelling.
@@
Together, these results highlight the dual scientific potential of PTA observations: as sensitive detectors of solar plasma phenomena and as monitors of intrinsic pulsar magnetospheric variability. The methodology developed here – combining detection of scientific DM-outliers, analyzing solar event data, and frequency-resolved profile analysis – can be systematically extended to the full InPTA-DR2 sample to uncover additional transient phenomena.
Future work will focus on (i) expanding this framework to other InPTA-DR2 pulsars with low ecliptic latitudes to identify further CME-related DM excursions, (ii) incorporating polarization and higher time-resolution data to explore the rotation variations due to solar events as well as physical triggers of mode changes, (iii) correlating DM and profile variability, and (iv) incorporating these propagation and intrinsic effects into the timing and noise modelling for PTA science. These efforts will deepen our understanding of both the heliospheric plasma environment and the dynamic magnetospheres of millisecond pulsars, thereby improving PTA data fidelity for gravitational-wave detection.
Acknowledgments

### v2

@@ 3.3 In situ data
The in situ data from STEREO-A reveals that the cme1, after starting from the Sun at 06-08-2022 01:48, with a speed of 900 km/s reaches STEREO-A around 08-08-2022 00:00 UTC; we see an enhancement around that time in the measured speed and temperature of the solar wind (see Fig. 4 - generated using CDAWeb [^7] Coordinated Data Analysis Web (CDAWeb) is part of NASA’s Space Physics Data Facility (SPDF), which serves as a central archive and access point for heliophysics and space weather data from various NASA and international missions - https://cdaweb.gsfc.nasa.gov/).
@@
The data coverage is incomplete for unknown reasons, which may explain the absence of the expected plasma-density jump. Nevertheless, we see subsequent increases in the plasma density after 10-08-2022 00:00 UTC, which notably corresponds to the cme2, which after starting from the Sun at 07-08-2022 11:00 UTC, with a speed of 776 km/s is supposed to reach STEREO-A around 09-08-2022 17:00 UTC. The differences between the predicted and observed arrival times of the CME likely arises from simplistic assumptions, such as constant CME velocity, which is not completely valid given the fact that CME can accelerate or decelerate along its path due to the drag force (CME_SW_drag_Maloney_1; CME_SW_drag_2; CME_SW_drag_3). The second CME (cme2) seems to be a possible secondary CME coming from nearly the same active region (as cme1, see Table 1), that got accelerated by possible interaction with the fast solar wind component.
@@ 3.4 DM excess estimate
Here, we estimate the DM excess from the in situ data and the spherically symmetric SW model. From a spherically symmetric SW model consideration, the formula for the DM contribution coming from a portion of the LOS, say $P_{1}$-$P_{2}$ (see Fig. 5(a)) is
$DM_{P_{1}-P_{2}}=\frac{n_{sat}(r_{sat}(\text{AU}))^{2}}{1~\text{AU}}\int_{\theta 1}^{\theta 2}d\theta,$ (1)
where $n_{sat}$ is the electron density at the position of the satellite of interest, $r_{sat}~(\text{AU})$ is the distance of the satellite from the Sun in units of AU, $\theta_{i}$ is the angle subtended by the sun-$P_{i}$ line on the earth-pulsar LOS, where $i\in\{1,2\}$.
@@
Using the position coordinates of the satellite of interest, the pulsar’s LOS, and the properties of the CME modelling as per DONKI, we identify the portion of the LOS that gets influenced by the cme2’s impact on the day of our CME-affected-observation (see Fig. 5(b)).
The value of $n_{sat}$ corresponding to cme2, as measured from the in situ data of STEREO-A, is roughly 10 cm-3. Integrating along the affected portion of the LOS, in Eq. 1, we get the DM due to both the CME and the background SW $\sim$ $5.7\times 10^{-3}$ pc cm-3. From the in situ data over a long time baseline, the background electron density on average is $\sim$ 5 cm-3. From this, the DM excess expected due to the background SW is $2.8\times 10^{-3}$ pc cm-3. Hence the DM excess from the in situ data is $\sim$ $2.9\times 10^{-3}$ pc/cc. This value conforms with the excess DM observed for PSR J1022+1001 on 09 August 2022 (see Section 3.1).
@@ 4 Analyzing a possible mode change epoch of PSR J2145-0750
[…]
Figure 6: DM time series of PSR J2145-0750 along the InPTA-DR2 baseline (all prominent outliers, which are artifacts of poor data, are not included in this plot), highlighting the single outlier, corresponding to a potential mode-change, with red circle. The blue points with vertical bars represent the DMCalc-evaluated DM with its uncertainty.
Interestingly, all of these small-time-scale deviations also correspond to good detections with high S/N, indicating the possibility of profile showing marginal variability w.r.t. its template. To examine further, the physical significance of the prominent outlier, and the small-time-scale deviations, we proceeded as follows.
@@ 4.1 Peak-ratio distribution
For each observational epoch, we updated the corresponding archive/fits files using the InPTA-DR2 final timing solution. The header DM values were then corrected with epoch-wise estimates obtained from DMCalc, and the profiles were subsequently dedispersed to remove dispersion-induced artifacts. For ready comparison and consistent analysis of all the profiles across the baseline, we positioned them at the same pulse-phase (kept the main peaks at the center). In addition, the profiles were band-equalized (Rana25) so that the power remained uniform across the frequency channels.
We then quantified the mode of the pulse profile by evaluating the peak ratio, defined as the ratio of the left peak to the right peak in the two prominent components of the profile, for each epoch. The resulting distribution of the ratio of the component amplitudes is well described by a Gaussian with mean and mode value of 1.0698 and a standard deviation of 0.0486, yielding 1-$\sigma$ bounds of [1.0212, 1.1183] (see Fig. 7).
@@
However, we identify a singular ‘outlier’, lying outside the Gaussian envelope, with a peak ratio of $\sim$0.8, observed at MJD 59468 (11 September 2021), the same outlier seen in the DM time series plot in Fig. 6, indicative of a possible mode change.
@@ 4.3 Eliminating systematics degenerate with mode change
The peak-ratio distribution reveals a single stable mode, i.e. the normal mode, well described by a Gaussian, and a distinct outlier corresponding to the epoch MJD 59468. The latter lies outside the Gaussian envelope and therefore raises the natural question of whether it corresponds to an abnormal mode, representing a genuine mode-change event or merely an artifact arising from propagation effects or observational systematics. Potential sources of such artifacts include: inter-stellar dispersion, which can alter the pulse shape; inter-stellar scintillation, which modulates the pulsar signal across frequency channels and can affect the frequency-integrated profile (particularly relevant for J2145–0750, which exhibits strong frequency-dependent profile variations); polarization calibration errors or flips in polarization channels; and radiometer noise, which may distort the band shape or modify the root-mean-square (RMS) of the profile baseline.
Each of these systematics has been carefully accounted for as follows. Dispersion effects are eliminated by using highly precise, epoch-wise DM estimates obtained with DMCalc on dual-band observations. Scintillation is accounted for by inspecting subbanded profiles for the four epochs of interest, which averages out the variation in signal strength across frequency channels constituting each of the subbands. Band-shape equalization and phase bin collapsing suppress the impact of radiometer noise by reducing the RMS of the noise floor. The polarization calibration related systematics would be ineffective in our dataset since they are total intensity data rather than the full four Stokes parameter recordings. It is for this reason, however, an intrinsic profile change in polarized components would remain undetected in our data. Additionally, PSR J2145-0750 was at a separation angle of $\sim 160^{\circ}$ with the Sun, at the potential mode-changing epoch MJD 59468, thereby ruling out any solar origin for this outlier. Furthermore, we checked the observatory settings and the observatory logs for that epoch MJD 59468 – we found no mention of any observatory related glitches or atmospheric events, that could have attributed to such a potential mode change; also the observatory settings for that day was the standard one used for InPTA observations.
@@ 4.4 Indication of potential intrinsic magnetospheric changes
Apart from the change in the relative heights of the peaks of the profile we also see a symmetric broadening of the pulse width. We explore this, as a final test, to establish the plausible theoretical reason behind the potential mode change. We convolve each of the subbands of the template with a symmetric Gaussian kernel and find the best fit (using $\chi^{2}$ test) which resembles the actual subbanded profile. The parameters used for this Gaussian convolution are $\{A,b,s,\sigma_{g}\}$, where $A$ is the overall amplitude scaling, $b$ is the base line offset, $s$ is the phase-shift, and $\sigma_{g}$ is the standard deviation. The value of $\sigma_{g}$ corresponding to the best-fit Gaussian kernel, which we denote by $\sigma$ [^8] This $\sigma$ is not to be confused with the one described in Section 4.1, provides an estimate of the observed broadening in the corresponding subbanded profile. From the results of this analysis (see Fig. 10), we find non-zero $\sigma$ (in bin units) for each of the subbands.
@@
(b) Subband 2 (central frequency = 461 MHz)
(c) Subband 3 (central frequency = 436 MHz)
(d) Subband 4 (central frequency = 411 MHz)
(e) Subband 5 (central frequency = 386 MHz)
(f) Subband 6 (central frequency = 361 MHz)
(g) Subband 7 (central frequency = 336 MHz)
(h) Subband 8 (central frequency = 311 MHz)
@@
$\Delta t_{\Delta DM}~(ms)=8.3\times 10^{6}\times\Delta DM\times\frac{\Delta\nu}{\nu^{3}}~=~\sigma~(ms)=\sigma~(bins)\times\frac{P}{nbin}~,$ (2)
where $P$ is the pulsar spin period, and $\Delta\nu$ is the width of the subband. We observe that the estimated $\Delta DM\sim 0.08~pc/cc$, for all the subbands, exceeds the DM uncertainty $\sim 0.001~pc/cc$, by nearly two orders of magnitude. Hence, such a huge error in DM would have been clearly visible as a parabolic drift in the frequency-phase plot of the dedispersed frequency-integrated profile, which we do not see. Therefore, the source for the observed pulse-width broadening is less likely to be due to dispersion effects. Scattering effect, if any, is also possibly not the dominating contributor to such a pulse-width broadening, as validated from the absence of well-defined asymmetric exponential tails of the individual pulses in the profile and that clear scintles of 3–4 MHz wide scintillation bandwidths were seen in the observations. The symmetric broadening in the pulse-widths, therefore, in all possibilities, indicate a potential magnetospheric origin.
5 Distinguishing CME-induced and mode-change-induced outliers
The two DM outliers investigated in this work arise from physically distinct mechanisms, and consequently required different criteria for their identification and establishing their physical origins. In this section we integrate the fundamental elements of the two case studies, explicitly comparing them for coherence.
In PSR J1022+1001, the identified outlier (MJD 59800) occurs at a small solar elongation and shows a clear DM excess above the modelled solar-wind contribution, with no accompanying change in the pulse shape – the criteria of small solar elongation is essential because heliospheric plasma structures can measurably affect the dispersion measure only when the line of sight passes reasonably close to the Sun. Since CMEs influence only the integrated electron density and do not alter the intrinsic properties of the pulsar, the absence of any profile-shape deviation is an essential discriminant that rules out intrinsic variability. The DM excess is subsequently found to be consistent with a CME intersecting the line of sight of the pulsar, as supported by DONKI WEC simulations, LASCO/C3 coronagraph imaging, and STEREO-A in situ plasma data.
For PSR J2145-0750, the situation is reversed. The identified outlier (MJD 59468) does not coincide with small solar elongation, but instead shows significant changes in pulse shape quantified by peak-ratio value well outside the stable distribution. Such morphology-altering behaviour at large solar elongation is indicative of a possible mode-changing event and cannot be produced by heliospheric propagation effects, as the pulsar line of sight passes far from the solar corona, rendering any CME-induced DM fluctuation negligible. This is further supported by symmetric pulse-width broadening across frequency subbands and rigorous checks against dispersion errors, scintillation, radiometer noise, and instrumental effects.
These two case studies demonstrate that DM outliers in PTA datasets arise from distinct physical mechanisms and therefore require different diagnostic criteria. CME-induced outliers manifest as excess DM at small solar elongations without accompanying profile changes, whereas mode-change events are characterized by simultaneous DM excursions and pulse-shape anomalies.
@@ 6 Conclusions
In this work, we investigated astrophysically motivated outliers in the InPTA-DR2 DM time series for two representative pulsars, PSR J1022+1001 and PSR J2145-0750. A key aspect of our approach is that the identification of meaningful outliers in PTA datasets cannot be based solely on statistical deviation but must be guided by physical context, since different mechanisms – such as heliospheric plasma structures or intrinsic mode change – imprint fundamentally different observational signatures. This context-driven framework naturally leads to distinct criteria for selecting outlier epochs in the two pulsars and provides a coherent interpretation for why their observational characteristics differ.
For PSR J1022+1001, we identified a distinct DM excursion, in comparison to the spherically symmetric solar-wind model prediction, on MJD 59800 (09 August 2022), which is geometrically consistent with a CME event detected by LASCO/C3 coronagraph, STEREO-A satellite and the DONKI database. Quantitative comparison between the DM excess inferred from radio data and that derived from in situ solar plasma measurements revealed excellent agreement, strongly establishing one of the first direct radio signatures of a CME along this particular PTA pulsar’s line of sight. This result underscores the potential for tracing heliospheric plasma dynamics and properties from the PTA datasets at astronomical distances where satellites are not present. Additionally, the number of IPTA pulsars monitored all across the sky is large enough, compared to space-based satellites, to possibly generate a three dimensional understanding of the outflow and interactions of solar wind and CMEs. The combination of several PTA datasets, across larger baselines, has the potential to contain more such solar-activity-induced outliers - this present understanding will therefore help in better modelling of the red-noise in the PTA dataset, induced by solar activities. This will lead to more accurate timing without having to throw away a significant portion of the dataset, at low separation angles, in the absence of better solar wind modelling.
@@
Together, these results highlight the dual scientific potential of PTA observations: as sensitive detectors of solar plasma phenomena and as monitors of intrinsic pulsar magnetospheric variability. The methodology developed here – combining detection of scientific DM-outliers, analyzing solar event data, and frequency-resolved profile analysis – can be systematically extended to the full InPTA-DR2 sample to uncover additional transient phenomena. Future work will focus on (i) expanding this framework to other InPTA-DR2 pulsars with low ecliptic latitudes to identify further CME-related DM excursions, (ii) incorporating polarization and higher time-resolution data to explore the rotation variations due to solar events as well as physical triggers of mode changes, (iii) correlating DM and profile variability, and (iv) incorporating these propagation and intrinsic effects into the timing and noise modelling for PTA science. These efforts will deepen our understanding of both the heliospheric plasma environment and the dynamic magnetospheres of millisecond pulsars, thereby improving PTA data fidelity for gravitational-wave detection.
Acknowledgements